OpenStudy (anonymous):

The mean of four consecutive even numbers is 15. The greatest of these numbers is ____ The least of these numbers is _____ please just tell me the answer

OpenStudy (anonymous):

well first lets find what the sum of the numbers could be by multiplying 15*4 since there are 4 number 15*4=60 so the sum of the 4 numbers =60

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

well mean=average and average is calculated by adding all of the numbers and dividing by the number of numbers added

OpenStudy (anonymous):

The greatest of these numbers is ____ The least of these numbers is _____

OpenStudy (anonymous):

the problem is not chain of consecutive numbers have the total of 60 13+14+15+16=58 and14+15+16+17=62

OpenStudy (anonymous):

so its .. The greatest of these numbers is 62 The least of these numbers is 58?

OpenStudy (anonymous):

no there is no chain of 4 numbers that adds up to 60 to create a mean of 15

OpenStudy (anonymous):

the greatest would have been 16 with the least of 13 but that still doesnt give us 60

OpenStudy (anonymous):

wait hold on even numbers

OpenStudy (anonymous):

sorry let me try again

OpenStudy (anonymous):

here we go ok so the chain of numbers would be 12, 14, 16 , and 18 18 is the greatest and 12 is the lowest

OpenStudy (anonymous):

is it 16 and 13???

OpenStudy (anonymous):

sorry misread it a bit

OpenStudy (anonymous):

18 and 12 now?

OpenStudy (anonymous):

yes I am deeply sorry

OpenStudy (anonymous):

its okay :) thank you for answering!

OpenStudy (anonymous):

no problem